OpenStudy (prince1342):

A circle has its center at (−2, −3) and a radius of 3 units. What is the equation of the circle?

OpenStudy (prince1342):

(x + 2)2 + (y − 3)2 = 9 (x − 2)2 + (y − 3)2 = 3 (x + 2)2 + (y + 3)2 = 9 (x + 2)2 + (y + 3)2 = 3

OpenStudy (prince1342):

@ganeshie8 Help please?

OpenStudy (anonymous):

Equation of a circle with center (h, k) and radius r is (x−h)2+(y−k)2=r2

rvc (rvc):

yup so substitute center(h,k)=(-2,-3)

OpenStudy (prince1342):

How did you get the formula?

OpenStudy (anonymous):

i looked it up @Prince1342